Part 1: Friday Line, a modeling agency, completed the following transactions during the first month of...

Part 1:

Friday Line, a modeling agency, completed the following transactions during the first month of operations:

April 1: Stockholders invested $65,000 cash along with equipment valued at $32,000 in the company in exchange for 1,000 shares of stock.

April 3: Friday Line paid $9,500 cash to settle accounts payable.

April 6: Friday Line completed services for a client and immediately received $2,500 cash.

April 9: Friday Line completed a $7,500 project for a client who must pay within 30 days.

April 12: Friday Line made credit purchases of $9,200 for office equipment and $4,100 for office supplies. Payment is due within 10 days.

April 19: Friday Line paid $3,200 cash for the premium on a 12-month insurance policy.

April 21: Friday Line pre-paid $12,000 for 12 months’ rent for their office space.

April 22: Friday Line received $6,000 cash as a partial payment for the work completed on April 9.

April 25: Friday Line purchased $725 of additional office supplies on credit.

April 26: Friday Line completed work for another client for $3,125 on credit.

April 27: Friday Line paid a dividend of $6,500 cash to its stockholders.

April 30: Friday Line paid $525 cash for this month’s utility bill.

Instructions:

Prepare journals for the above economic transactions and include the effect that each transaction has on the accounting equation. Use the following assignment template for Friday Line Co.

Part 2

The unadjusted trial balance of Bolton Consulting is entered on the partial worksheet below.

Bolton Consulting

Work Sheet

For the year ended December 31



Account


Unadjusted Trial Balance



Adjustments


Adjusted Trial Balance


Income Statement

Balance Sheet and Statement of Stockholders’ Equity


Debit

Credit

Debit

Credit

Debit

Credit

Debit

Credit

Debit

Credit

Cash

31,000










Accounts Receivable

8,000










Supplies

12,500










Automobiles

175,000










Accum. Depr. - Autos


52,000









Accounts payable


17,500









Unearned fees


26,500









Salaries payable











Capital Stock


58,000









Dividends

42,500










Fees earned


241,750









Salary expense

91,000










Rent expense

35,750










Advertising expense











Supplies expense











Depreciation expense

______

______









Totals

395,750

395,750
Answer #1
Part 1)
Journal Entries: Effect on Accounting Equation:
Date Acc Title Debit $ Credit $ Assets: Liabilities+ Share Equity
1-Apr Cash 65000 Cash other Current Assets Fixed Assets = Liabilities + common stock + Ret. Earnings
Equipment 32000 65000 32000 97000
Common stock 97000
(being capital infused in company - 1000 shares)
3-Apr Accounts payable 9500 -9500 = -9500
Cash 9500
(creditors paid)
6-Apr Cash 2500 2500 = 2500
Service revenue 2500
(services revenue cashed)
9-Apr Accounts receivable 7500 7500 = 7500
Service revenue 7500
(being service revenue recorded)
12-Apr Office Equipment 9200 4100 9200 = 13300
Office supplies 4100
Accounts payable 13300
(office equipment and supplies purchased on credit)
19-Apr Prepaid Insurance 3200 -3200 3200 =
Cash 3200
(insurance purchased for a year)
21-Apr Prepaid rent 12000 -12000 12000 =
Cash 12000
(rent paid for a year)
22-Apr Cash 6000 6000 -6000 =
Accounts receivable 6000
(collection of debtors)
25-Apr Office supplies 725 725 = 725
Accounts payable 725
(supplies purchased on credit)
26-Apr Accounts receivable 3125 3125 = 3125
Service revenue 3125
(being service revenue recorded)
27-Apr Dividend 6500 -6500 = -6500
Cash 6500
(being dividend paid to shareholders)
30-Apr Utilities Expenses 525 -525 = -525
Cash 525
(utilities expenses incurred) 41775 24650 41200 = 4525 97000 6100
107625 = 107625

Part II) The things are not clear and figures are scattered please, so can't reply.